OpenStudy (anonymous):

solve the following system show your work 3x-2y+2z=30 -x+3y-4z=-33 2x-4y+3z=42 I really don't know how to do this if someone could guide me through i would be so thankful

OpenStudy (amistre64):

matrix method is usually simplest for this, if you have a calculator

OpenStudy (anonymous):

how would i do it?

OpenStudy (amistre64):

if we do substitution, then solve the first equation for a variable, and sub it into the variable of the other 2

OpenStudy (amistre64):

can you do substitution?

OpenStudy (anonymous):

i'm not very good can i get an example?

OpenStudy (amistre64):

ax + by + cz = k solve for z: subtract off the non z terms, and divide off the c

OpenStudy (amistre64):

then for some setup: mx + ny + pz = j we replace z with its equivalent xy setup

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

so add xy then divide out c?

OpenStudy (amistre64):

the x and y terms so that z is defined in terms of x and y

OpenStudy (anonymous):

ok so for this problem i start with the first equation and do that?

OpenStudy (amistre64):

we can yes

OpenStudy (anonymous):

or should we Choose the easiest way to eliminate a variable, and then write a system of two equations

OpenStudy (amistre64):

if you like elimination method then we can do that as well. there are a few different approaches we can take, they all give the same end results

OpenStudy (anonymous):

ok i think i know that one the linear combination method?

OpenStudy (amistre64):

use eq1 to eliminate the x parts of the last 2 3x-2y+2z=30 -x+3y-4z=-33 < times by 3 2x-4y+3z=42 < times by -3/2

OpenStudy (amistre64):

i not sure what a linear comination method is, unless thats another way of saying matrix solution

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

ok so i would times the whole equation my 3in order to get the y eliminated for the second euation

OpenStudy (anonymous):

if so the second equation is 15x+15z=216 right?

OpenStudy (amistre64):

3x-2y+2z=30 -x+3y-4z=-33 < times by 3 2x-4y+3z=42 < times by -3/2 3x-2y+2z=30 -3x+9y-12z=-99 -3x+6y-9z/2=-63 3x-2y+2z=30 0x+7y-10z=-69 0x+4y-5z/2=-33 7y-10z=-69 8y- 5z=-66 now its a 2 variable setup

OpenStudy (anonymous):

oh ok that makes sense

OpenStudy (amistre64):

7y-10z=-69 8y- 5z=-66 << times -2 7y-10z=-69 -16y+10z=132 -------------- -9y =63 y = -7

OpenStudy (amistre64):

using eq2 we find 7(-7)-10z = -69, solve for z knowing y and z, solve for x in the first equation

OpenStudy (anonymous):

OK thank you i appreciate your help!

OpenStudy (amistre64):

youre welcome, when you find a solution, let me know what you get.

OpenStudy (anonymous):

ok i will one moment

OpenStudy (anonymous):

im stuck so on the first equation the ending equation was 2x-2z=-3 was that right?

OpenStudy (amistre64):

we know y = -7, what did you get for z?

OpenStudy (anonymous):

i added the first and second equation

OpenStudy (amistre64):

i add the 1st and 2nd, the 1st and 3rd and got rid of x 3x-2y+2z=30 -x+3y-4z=-33 < times by 3 2x-4y+3z=42 < times by -3/2 3x-2y+2z=30 -3x+9y-12z=-99 -3x+6y-9z/2=-63 3x-2y+2z=30 0x+7y-10z=-69 0x+4y-5z/2=-33 7y-10z=-69 8y- 5z=-66 now its a 2 variable setup

OpenStudy (amistre64):

since we have a 10 and 5, i figure times 2 will get rid of z for us

OpenStudy (amistre64):

7y-10z=-69 8y- 5z=-66 << times -2 7y-10z=-69 -16y+10z=132 -------------- -9y =63 y = -7

OpenStudy (amistre64):

can you get this far?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then using: 7y-10z = -69 we can find z what is z?

OpenStudy (anonymous):

z= 2?

OpenStudy (amistre64):

good, y = -7, z=2 now we can use any of the original equations .... if we use equation 1 3x-2y+2z=30 3x-2(-7)+2(2)=30 solve for x

OpenStudy (anonymous):

x=4?

OpenStudy (amistre64):

yes, good job

OpenStudy (anonymous):

thank you!

OpenStudy (amistre64):

yw